OpenStudy (anonymous):

for a circle of radius 9 feet, find the arc length subtended by a central angle of 52 degrees? A. 13pi/5 b. 11pi/5 c. 13pi/11 d. 5pi/11

OpenStudy (anonymous):

First; convert 52° into radians 52° / 180 = 13 / 45 pi (radian) s = radian x radius s = (13/45) x 9 s = 2.6 ft / pi x pi(3.14) = 8.16 ft

OpenStudy (anonymous):

@chlobohoe

OpenStudy (anonymous):

i got the answer i think thank you!!!

OpenStudy (anonymous):

Yup!